OpenStudy (anonymous):

Need more tutoring then answers so if your willing to teach me pllz help Medalls!!

OpenStudy (anonymous):

With what?

OpenStudy (anonymous):

with my algerbra

OpenStudy (anonymous):

I'm happy to help post your questions

OpenStudy (anonymous):

okay like i dont understand how to multiply fractions

OpenStudy (anonymous):

for example 3/2 x1/2

OpenStudy (unklerhaukus):

\[\frac ab\times\frac cd=\frac{a\times c}{b\times d}\]

OpenStudy (abb0t):

\[\frac{ 3 }{ 2 } \times \frac{ 1}{ 2 } = \frac{ 3 \times 1 }{ 2 \times 2 }\]

OpenStudy (zugzwang):

^What he said... I actually find multiplying fractions a lot less tricky than adding or subtracting them.

OpenStudy (anonymous):

And if you invert it then you are dividing :D

OpenStudy (zugzwang):

Provided you have mastered reducing fractions to lowest terms.

OpenStudy (anonymous):

*the second fraction

OpenStudy (anonymous):

okay so the answer should be 3/6

OpenStudy (anonymous):

Mmm this is true

OpenStudy (zugzwang):

Well, it must be 3/6... But you can do better than that~ Lowest terms...?

OpenStudy (anonymous):

1/2

OpenStudy (zugzwang):

Wait a sec. what was the question to which your answer was 3/6 ?

OpenStudy (anonymous):

wait so any time the number is odd i the lowest term always gonna be 1

OpenStudy (anonymous):

3/2 x 1/3

OpenStudy (anonymous):

my answer was 3/6 then you said lowest terms so i said 1/2 was I right

OpenStudy (zugzwang):

Yes. Let's have another one: \[\frac{4}{3}\times\frac{9}{4}\]

OpenStudy (anonymous):

36/12

OpenStudy (zugzwang):

Yes... continue...

OpenStudy (anonymous):

but how would i continue if both of there lowest terms are 2

OpenStudy (zugzwang):

2? How do you figure 2?

OpenStudy (anonymous):

well if you do 36>18>9>3>1 12>6>>2 oh i see

OpenStudy (anonymous):

so it be 1/2 right

OpenStudy (zugzwang):

NO... Remember that in the end, fractions are just compact division: \[\large \frac{a}{b}=a \div b\] So what is 36 divided by 12?

OpenStudy (anonymous):

3

OpenStudy (zugzwang):

And there is your answer.

OpenStudy (anonymous):

so the answer is flat out 3

OpenStudy (zugzwang):

Yes. Remember that this can happen. Anyway, a whole number is technically a fraction with a denominator of 1, in lowest terms.

OpenStudy (anonymous):

so if i wanted to make it into fraction form id just 3/1

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

yeah